OpenStudy (anonymous):

A cylindrical cardboard mailing tube has a diameter of 3 inches and a height of 28 inches a total of 23% of the tubes capacity is filled with spherical packing foam. Each foam sphere has a diameter of 0.8 inches. Approximately how many pieces of the foam are in the tube? Use 3.14 to approximate pi and round off your answer to the nearest whole piece of foam.

OpenStudy (radar):

Determine volume of the cylinder Determine what 23% of this volume is (make a note of that value) Next determine volume of a foam sphere. Divide the noted volume by the volume of foam sphere Round this to the nearest whole number.

OpenStudy (radar):

Here are some equations you may need. Volume of cylinder V=Pi*r^2*h r= radius, h=height. Volume of sphere V=(4/3)*Pi*r^3